OpenStudy (anonymous):

What is the y-intercept of the function? y=20,000(0.95)^x A) 20,000 B)0.95 C)1 D)0

OpenStudy (lgbasallote):

change x to 0 then solve for y

OpenStudy (unklerhaukus):

\[y(x)=20,000(0.95)^x\]the y intercept \((y_0,0)\) is when x=0 \[y_o=y(0)=20,000(0.95)^0=\] use \[n^0=1\]